OpenStudy (anonymous):

factor completely 5x^10-80

OpenStudy (anonymous):

factor out the 5 first and start with \[5(x^{10}-16)\]

OpenStudy (anonymous):

next step, factor \(x^{10}-16\) as the difference of two squares \[a^2-b^2=(a+b)(a-b)\] with \(a=x^5,b=4\)

OpenStudy (anonymous):

those are the steps